# A Benchmark for Optimal Multi-Modal Multi-Robot Multi-Goal Path Planning with Given Robot Assignment

### v1

@@ IV-2 Locally informed sampling
In our setting, the optimal paths are typically long compared to the size of the state space, and therefore, the informed set is typically larger than the valid configuration space (similar as in [35]).
@@
Given that we do know that a solution needs to pass through some regions in space, the standard approach of taking the cost function as $c_{\text{lb}}$ is not approximating the true path cost well. The main problem then is that the path constraints from the tasks are constraints for subspaces of the full composite space (i.e., even if we know that robot $r_{i}$ needs to be at pose $q_{g_{1}}$ first, robot $r_{j}$ might be unconstrained), and it is thus not trivial to construct a good lower bound for the path-cost. Additionally, compared to the euclidean norm, our cost function does not easily allow for direct sampling from the informed set, therefore requiring rejection sampling, which can be compute intensive.
@@
To alleviate these issues, we propose locally informed sampling: instead of using the start and the goal pose in Eq. 4, we randomly sample two points $q_{i},q_{j}$ on the path, and do informed sampling with these points:
$\mathcal{Q}_{\text{lis}}=\{q\,|\,c(q_{i},q)+c(q,q_{j})\leq c^{\pi}(q_{i},q_{j})\},$ (5)
@@
where $c^{\pi}(\cdot)$ is the cost to reach a point on the path along the best found path. We extended this to the sampling of modes by sampling from the set of modes that are in-between the modes that $q_{i},q_{j}$ are in respectively. This set of in-between modes can be inferred from the task dependency graph.
1 $T\leftarrow\{q_{0}\}$, $M\leftarrow\{m_{0}\}$
2 while not terminated do
3     $q_{\text{rnd}},m_{\text{rnd}}\leftarrow$SampleWithGoalBias()
4     $n_{\text{close}}\leftarrow$T.get_nearest($m_{rnd}$, $q_{rnd}$)
5     $n_{\text{new}}\leftarrow$Steer($n_{\text{close}},q_{\text{rnd}}$)
6     T.add($n_{\text{new}}$)
7     T.rewire($n_{\text{new}}$)
8     if $n_{\text{new}}$ is subgoal
9          $M\xleftarrow{\scriptscriptstyle+}\texttt{getNextMode($n_{\text{new}}$)}$
10
11    if improved solution and do shortcutting
[…]
13          T.addPath(sol)
14
Algorithm 1 Composite Multi Goal RRT*.
[…]
2 while not terminated do
3     $G\leftarrow$SampleTransitionNodes()
4     $G\leftarrow$SampleReachedModes()
5     if GoalNodeFound()
[…]
7          if improved solution and do shortcutting
[…]
9               G.addPath(sol)
10
11
return sol
Algorithm 2 Composite Multi Goal PRM*.
1 while not terminated do
2     $r\leftarrow$sampleRobot()
3     $i,j\leftarrow$sampleIndices()
4     if IsShortcuttable($r,i,j$)
5          $p\leftarrow$constructShortcutPath(sol, $r,i,j$)
6          if couldImproveCost($p$) and isCollisionFree($p$)
[…]
8
9
10
return sol
Algorithm 3 Multi-robot multi-modal partial shortcutting.
IV-A Composite RRT*
We introduce the required changes to a standard RRT* planner [19, 36] to make it work in the multi-modal multi-goal setting (Algorithm 1): The core ideas are the same - we randomly sample a state consisting of a pose and a mode. We maintain a set of previously reached modes $M$, which is initialized with the start mode. We then sample a valid mode/pose combination either by sampling a mode uniformly at random from the reached modes $M$, and then sampling a pose uniformly in this mode, or if we previously found a path, we use locally informed sampling to sample the mode and the pose. We then search the closest state in our tree (according to the distance function), and grow the tree towards the randomly sampled state. The new state is added to the tree and the tree is rewired. If the new state is a transition (i.e., a pose fulfilling the goal constraints of a task), we add the mode that we reach from this transition to the set of reached modes. We do shortcutting (Section IV-C) once a solution is found to improve convergence of the planners [37].
@@ IV-A Composite RRT*
We also introduce a bidirectional variant, which samples goals in each mode (which are used as starts in the next mode), and then runs bidirectional RRT* in each mode.
@@
IV-B Composite PRM*
We introduce the required changes to a PRM-based planner [20] to adapt it to our setting (Algorithm 2). We build our roadmap by first sampling a batch of transition nodes (i.e., nodes that fulfill the constraints of a task), which enlarges the set of reached modes. We then sample a batch of nodes in the modes that we reached so far (using locally informed sampling if we found a solution, uniformly at random if not). If we found a transition that allows us to reach to goal, we then run an A* search over the multi-modal graph formed by all the samples so far (which is similar to MLA* [11]). We implement this search using an edge queue, which allows delaying edge validation until the edge is inspected, therefore minimizing collision checking effort. To select neighbors, we choose a radius based approach, with the radius determined as in [36]. Similar to the RRT planners, we do shortcutting after a path was found that improves the current solution.
@@ IV-B Composite PRM*
For the search of the graph that we are incrementally building, we require a heuristic. The heuristics in our setting present a tradeoff between being useful and being fast to compute. We compute a heuristic by running a reverse search on the transition nodes to compute a lower bound on the cost to reach a mode. The heuristic is then the sum of the cost to reach a specific transition and the cost to reach the goal from this transition mode: $h(q)=\min_{i}c(q,q^{\text{trans}}_{i})+c_{lb}(q^{\text{trans}}_{i})$, where $q^{\text{trans}}$ are the poses belonging to transition nodes in a mode.
IV-C Postprocessing
@@
We also introduce a path post-processing method that is specifically tailored to the multi-goal motion planning problem. Our approach (Algorithm 3) builds on top of partial shortcutting [38]. The main idea is that we do not shortcut all dimensions of the path at once, but select a subset of dimensions from the composite space, and try to shortcut only this subset, while keeping the rest of the plan the same. Instead of uniformly sampling from all possible subsets, we notice that in practice it is more efficient to sample a robot, and then shortcut all robot dimensions at once. This enables improving the unconstrained parts of the path (i.e., the parts that do not influence a mode-transition), while still keeping the simplicity of the shortcutting approach.
IV-D Proof sketches
@@
Both planners uniformly sample all transitions, and sample uniformly within the modes, therefore eventually discovering all possible modes. Within the modes, both planners work as their respective base version, and are therefore probabilistically complete, and will eventually connect to the transitions. If a path exists, the discovered transitions will lead to the terminal mode, which contains the goal pose. Thus, both planners will find a path if one exists, and are therefore probabilistically complete.
@@
The proof sketch for almost-sure asymptotic optimality follows similar logic: The A* search that we run in the PRM-planner is optimal, and therefore, as the number of uniform-random samples approaches infinity, we converge to the optimal paths within the modes, and to the optimal mode-transitions. Similarly, the rewiring-step in the RRT-planners is the same as in RRT*, and thus inherits the optimality guarantees from RRT* even over multiple modes.
V Benchmark & Experiments
(a) 2D hallway.
(b) 2D random.
(c) 4-arm box-stacking.
(d) Mobile assembly.
RRT*BiRRT*PRM*
Figure 5: Evolution of median cost over time along with the 95% non-parametric confidence intervals over 50 runs.
@@ V Benchmark & Experiments
The main contribution of this work is the open source benchmark implementing a variety of scenarios with a diverse sets of robots and tasks. The 21 base scenarios range from simple settings where the optimal solution path is known to help validate properties of the planner, to more complex scenarios with up to 5 robots (with a total of 30 degrees of freedom), and up to 22 goals. For many of the problems, there is both a version with a full task ordering, and one using a dependency graph as task specification. We show a selection of scenarios in Fig. 4.
@@
As the main goal of this work is the introduction of the problem statement and making the problem accessible, and not the fastest possible planner, the benchmark and planners are implemented using Python, while the computationally expensive parts, i.e., collision checking and kinematics are leveraging the python bindings of RAI [^2] https://github.com/MarcToussaint/robotic. This backend can however easily be replaced, allowing for, e.g., GPU parallelization if the backend supports it.
@@ V-A Experiments
We present a selection of tasks to showcase the different types of scenarios and task sequence specifications and compare the planners on them. We report results for four representative scenarios: a 2D scenario with two robots that is similar to the classic wall-gap, where the robots have to switch positions and go back again (6D configuration, 3 subgoals, Fig. 4(a)), a 2D scenario with 3 agents that have to reach a number of goals in sequence (9D conf. space, 13 subgoals, Fig. 4(b)), a scenario with 4 robot arms where 8 boxes have to be placed on top of each other (28D, 17 subgoals, Fig. 1), and a scenario involving 4 mobile manipulators rearranging a wall (24D, 17 subgoals, Fig. 4(f)). In these scenarios, the task sequences are fully determined for the two 2D scenarios, and only partially given for the robot-scenarios, i.e. the order and timing of each task needs to be figured out by the planner. The tasks are assigned randomly to a robot, and in all these instances, there is a single goal pose for each subgoal.
If not specified otherwise, we consider the cost function with $w=0.01$, i.e., the min-makespan problem with a small path-length regularization, and we report the median solution costs over 50 runs along with confidence intervals.
The experiments were run with Python 3 on a Ryzen 7 5800X (8-core, 4’491 MHz) and 32 GB RAM.
@@ V-B Results
Figure 5 shows the resulting cost evolution plots. In general, it is expected that the solution times for these problems are higher than solution times for a ’standard’ path planning problem. This is partially due to the length of the plans, and therefore the larger collision checking effort, in addition to the large search space, and partially due to the problem structure: If the planner finds a path to a better mode-transition, the rest of the path has to be completely replanned for all following modes and mode-transitions.
The RRT-based planners are up to an order of magnitude faster at finding an initial solution, while the PRM-planner tends to converge faster to the optimal solution once an initial path was found. Of the RRT-planners, the bidirectional version is usually faster.
@@ V-C Ablations
In the following, we present ablations to explain design decisions of the planning algorithms.
@@ V-C1 Locally informed sampling
As alluded to earlier, only using naive sampling in the problems we consider here is insufficient, as the informed set is large compared to the configuration space, and thus in many cases does not help much even compared to only using uniform sampling. In Fig. 7, we show the planners using locally informed sampling, and compare it to the naive informed sampling. The planners using locally informed sampling outperform the planners that do not. However, it can also be seen that in the higher dimensional environment, the optimum is not reached in the allotted planning time by either of the groups of planners.
RRT*BiRRT*PRM*
Figure 7: Local (solid line) and global (dashed line) informed sampling on the 2D hallway (left) and the 4-arm box-stacking (right) scenarios. The dashed gray line is the best cost achieved by the non-ablated planners.
RRT*BiRRT*PRM*
Figure 8: Planners with (solid line) and without (dashed line) shortcutting on the 2D hallway (left) and the 4-arm box-stacking (right) scenarios. All planners are using locally informed sampling. The dashed gray line is the best cost achieved by the non-ablated planners.
@@ V-C2 Shortcutting in the planner
Shortcutting a path after it was found leads to faster convergence than using only informed sampling, as can be seen in Fig. 8. Shortcutting in the planner does not affect the optimality or completeness properties of the planners, but does significantly improve convergence by quickly cutting down on the maximum cost that the planners need to consider. This effect is more pronounced in high-dimensional scenarios.
V-C3 Connection strategy
Using k-nearest neighbors in the PRM-planner results in finding the initial solution faster, but slower convergence to the optimal solution (see Fig. 9). We find that this is the case since the samples that we have are non-uniformly distributed around the transitions. The difference in speed can largely attributed to the fact that many more connections are added to the queue in the radius based connection strategy compared to the k-nearest approach. This leads to a higher branching factor, and thus a slower planner.
PRM* (radius)PRM* (k-nearest)
Figure 9: Different connection strategies for the PRM planner on the 2D hallway (left) and the 4-arm box-stacking (right) scenario.
@@ VI Discussion & Limitations
There are some limitations on the scaling of the planners: that they do not scale well both with the number of robots (due to the increase in state-space dimensions) and they do not scale well with the number of goals. These limitations stem mostly from the decision to plan in the composite space and to treat the complete problem as one single big planning problem.
While we introduce a problem formulation that deals with dependency graphs, this increases the size decision space, and we believe that there could be a better way to approach the problem than a brute force search, which the sampling based planners do here.
Despite these limitations, we find that the planning in composite space is likely good enough for many complex industrial settings, where the number of used robots does not go much beyond four robots. We believe that a promising approach for improving the planner performance is leveraging the structure of the problem space, i.e., using the fact that we are dealing with multiple robots and multiple sub-goals, by, e.g., not planning over the full horizon directly, but only planning for a finite horizon.

### v2

@@ IV-3 Locally informed sampling
In our setting, the optimal paths are typically long compared to the size of the state space. Therefore, the informed set is typically larger than the valid configuration space (similar as in [37]), and thus not helpful.
@@
Given that we do know that a solution needs to pass through some regions in space (the goals that we need to reach), the standard approach of taking the cost Eq. 1 as $c_{\text{lb}}$ is not approximating the true path cost well. It is not easy to construct a good lower bound for the path-cost, since the regions that we need to pass through are subspaces of the full composite space (i.e., even if we know that robot $r_{i}$ needs to be at pose $q_{g_{1}}$ first, robot $r_{j}$ might be unconstrained, thus leading to a volume that we need to pass through instead of a single point). Additionally, compared to the Euclidean norm as cost, our cost function does not easily allow for direct sampling from the informed set.
@@
To alleviate these issues, we propose locally informed sampling: instead of using the start and the goal in Section IV-3, we randomly sample two points $q_{i},q_{j}$ on the previously found path, and do informed sampling with these points:
$\mathcal{Q}_{\text{lis}}=\{q\,|\,c(q_{i},q)+c(q,q_{j})\leq c^{\pi}(q_{i},q_{j})\},$
@@
where $c^{\pi}(\cdot)$ is the cost to reach a point on the path along the best found path. We extended this to the sampling of modes by sampling from the set of modes that are in-between the modes that $q_{i},q_{j}$ are in respectively.
IV-A Considerations for Multi-Modal, Multi-Goal Planners
We now discuss changes that deviate from the standard approach of the planners that we adapt.
IV-A1 Composite RRT*
Compared to standard RRT* [19, 33], we maintain a set of previously reached modes $M$ (i.e., a mode-graph similar to Fig. 3), which is initialized with the start mode. In order to sample configurations in the planner, we uniformly sample poses within the limits of the robots, and sample modes from this set. If we previously found a path, we use locally informed sampling to sample the mode and the pose. Further, we bias the random sampling in the RRT to the goal in a mode, instead of the terminal goal. Other than this, the standard versions of randomly sampling states and connecting to them is used.
@@ IV-A1 Composite RRT*
We also introduce a bidirectional variant (BiRRT*), which samples goals in each mode (which are used as starts in the next mode), and then runs bidirectional RRT* in each mode.
@@
IV-A2 Composite PRM*
The PRM* planner also follows the standard version closely. The main change is that the roadmap is constructed by first sampling a batch of transition nodes (i.e., nodes that fulfill the constraints of a task), which enlarges the set of reached modes and then sampling a batch of nodes in the reached modes, until a goal mode is reached. Then, a standard A* with an edge queue is used for the search over the roadmap.
@@ IV-A2 Composite PRM*
Since A* needs a heuristic in order to be effective, and the euclidean distance is not informative in our setting, we compute a heuristic for the A* search by running a reverse search on the transition nodes to compute a lower bound on the cost to reach a mode. The heuristic is then the sum of the cost to reach a specific transition and the cost to reach the goal from this transition mode:
$h(x=(q,m))=\min_{i}\left[c(q,q_{i}^{m,\text{trans}})+c_{\text{rev}}(q_{i}^{m,\text{trans}})\right]$ (2)
where $q^{m,\text{trans}}$ are the poses belonging to the transition nodes in a mode.
IV-A3 Composite IT-planners
AIT* and EIT* [38, 34, 39] use adaptive problem-specific heuristics to focus the search on promising regions. These heuristics are computed and refined through an asymmetric bidirectional search to obtain more informative exploration and improved solution quality.
Search
To explore the graph, AIT* and EIT* perform their respective asymmetric bidirectional search strategies, as described in [34]. Although each planner utilizes distinct heuristics, both follow an identical approach to handle transitions: When the forward search reaches a transition node, it only expands to valid successor modes, enforcing forward-directed connectivity; the reverse search applies the same logic in opposite direction (i.e., expanding towards previous modes). Both planners need an admissible a priori cost-to-come (i.e., from the start to the state $x$) heuristic - we use a modified version of Eq. 2 with the lower bound cost to reach a mode computed from the start node instead of the goal (i.e., replacing $c_{\text{rev}}$ with the cost computed by running a forward search only over the transition nodes of the previous modes). This heuristic is a trade-off between heuristic accuracy and computational efficiency.
The reverse search in EIT* incorporates heuristics approximating the computational effort required to find a valid path. We define the computational effort as the collision checking effort. Following this definition, the effort-to-come is computed analogously to the cost-to-come heuristic in Eq. 2, where cost terms are substituted by effort.
IV-B Postprocessing
@@
When a path is found in any of the planners, we run a postprocessing step to rapidly improve the path cost, which decreases the volume of the informed set. Our approach is a minor modification of partial shortcutting [40]: The main idea is that we do not shortcut all dimensions of the path at once, but select a subset of dimensions from the composite space, and try to shortcut only this subset, while keeping the rest of the plan the same. Instead of uniformly sampling from all possible subsets, we notice that in practice it is more efficient to choose a robot, and then shortcut all dimensions of this robot at once. This enables improving the unconstrained parts of the path (i.e., the parts that do not influence a mode-transition), while still keeping the simplicity of the shortcutting approach.
IV-C Proof sketches
@@
All planners uniformly sample all transitions, and sample uniformly within the modes, therefore eventually discovering all possible modes. Within the modes, both planners work as their respective base version, and are therefore probabilistically complete, and will eventually connect to the transitions. If a path exists, the discovered transitions will lead to the terminal mode, which contains the goal pose. Thus, both planners will find a path if one exists, and are therefore probabilistically complete.
@@
The proof sketch for almost-sure asymptotic optimality follows a similar logic: The search that we run in the PRM and EIT/AIT-planners is optimal, and therefore, as the number of uniform-random samples approaches infinity, we converge to the optimal paths within the modes, and to the optimal mode-transitions. Similarly, the rewiring-step in the RRT-planners is the same as in RRT*, and thus inherits the optimality guarantees from RRT* even over multiple modes.
V Experiments
(a) 2D hallway (*).
(b) 4-arm-box stacking (*).
(c) Mobile assembly (**).
(d) 2-arm rearrangement (***).
RRT*BiRRT*PRM*AIT*EIT*Prioritized
Figure 5: Evolution of median cost over time along with the 95% non-parametric confidence intervals over 50 runs. We also show the median initial solution time and cost using the square with error bars. (*) indicates a fully given task sequence, (**) indicates a partial task ordering, (***) indicates unassigned and unordered tasks. Note that the prioritized planner only acts as anytime planner and improves its cost in the (**) and (***) settings, where multiple different sequences can be generated and planned for.
@@ V Experiments
The main contribution of this work is the formulation of the multi-modal, multi-robot, multi-goal planning problem and the adaption of standard sampling-based planners to the formulation. We implement this formulation in a variety of base scenarios in an open-source benchmark featuring diverse sets of robots and tasks. The 21 base scenarios range from simple settings where the optimal solution path is known to help validate properties of the planner, to more complex scenarios with up to 5 robots (with a total of 30 degrees of freedom), and up to 74 goals. For many of the problems, there are versions with a complete task ordering and assignment, and others that are either unordered, partially ordered, or where tasks are unassigned. We show a selection of scenarios in Fig. 4.
@@
We provide proof-of-concept implementations of the planners in Python, while the computationally expensive parts, i.e., collision checking and forward kinematics use a more performant backend, which can be easily replaced, allowing for, e.g., GPU parallelization if the backend supports it. Currently, the benchmark supports implementations of environments in Pinocchio [41] and RAI [^3] https://github.com/MarcToussaint/robotic.
@@ V-A Experiments
We present a selection of experiments to showcase the different types of scenarios and task sequence specifications and compare the optimal planners to a prioritized planner [42] on them [^4] Synchronized planners and standard MAPF planners are not applicable to these problems, and are thus not part of the benchmark.. We report results for four representative scenarios: (a) A 2D scenario with two robots that is similar to the classic wall-gap, where the robots have to switch positions and go back to their start poses (6D conf. space, 3 subgoals, Fig. 4(a)), (b) a scenario with 4 robot arms where 8 boxes have to be stacked (24D conf. space, 17 subgoals, Fig. 1), (c) a scenario involving 4 mobile manipulators rearranging a wall (24D conf. space, 17 subgoals, Fig. 4(f)), and (d) a scenario with 2 robot arms that have to collaborate with handovers to rearrange and reorient 9 boxes (12D conf. space, 28 subgoals, Fig. 4(e)). In these scenarios, the task sequence is fully determined for scenario (1) and (2), partially given for scenario (3) and not assigned or ordered for scenario (4). To deal with partially ordered problems in the prioritized planner, we generate valid random sequences, and plan for those until the time runs out.
We use the cost function with $w=0.01$, i.e., the min-makespan problem with a small path-length regularization in scenarios (b)-(4d), and $w=1$ in scenario (a). We report the median solution costs over 50 runs with different random seeds. We use a frontier mode sampling strategy until a path is found, and then switch to uniform mode sampling in all planners if not stated otherwise.
The experiments were run with Python 3.10 on a Ryzen 7 5800X (8-core, 4’491 MHz) and 32 GB RAM.
Figure 6: The path at different iterations in the process, from the initial to the optimized solution. Color progression indicates time from start to end (rectangle , disk ) .
@@ V-B Results
Figure 5 shows the resulting cost evolution plots. Figure 6 shows the evolution of a path for the hallway problem.
Comparing the optimal planners introduced in this work and the prioritized planner shows that the suboptimal planner finds reasonable solutions (between 20% and 70% over the optimal solution, but much better than the initial solutions) quickly. As expected, the optimal planners converge to solutions with lower cost, but do so at a later time. Surprisingly, the planners that plan in composite space find initial solutions faster than the suboptimal planner in some settings. We explain this mainly with the fact that treating the previously planned paths as fixed sometimes constrains the planning space such that the planning problem becomes considerably harder compared to the setting where all agents can be moved around freely.
Convergence in complex problems
In scenarios (c) and (d) the planners do not reliably converge to the same solution at the end of their runtime but remain 5%-10% above the lowest found solution, showing that some of the found solutions are not optimal. This is because the problem space is extremely large due to the discrete choice of task assignment and ordering in addition to the high dimensionality of the continuous space. Thus, while the planners are theoretically optimal, in practice, we find that the planners tend to get stuck in a local optimum and do not switch the discrete assignment and ordering often. Note that in scenario (c), the median final cost of the prioritized planner is lower than the median final cost of some of the composite space planners. This is due to the better exploration of different sequences.
@@ V-C Ablations
We present ablations to explain the design decisions that deviate from the standard versions of the planners. In the following, we do not show the results for all planners as the effects are similar for most, and would crowd the figures.
@@ V-C1 Locally informed sampling
Only using standard informed sampling in the problems we consider here is insufficient, as the informed set is large compared to the configuration space, and thus effectively equivalent to uniform sampling. The planners do not use shortcutting in order to better highlight the difference between the ablations. The planners using locally informed sampling outperform the planners that do use standard informed sampling (Fig. 8).
BiRRT*PRM*EIT*
Figure 8: Local (solid) and global (dashed) informed sampling on the 2D hallway (left) and the 4-arm box-stacking (right) scenarios. The dashed blue line is the best cost of the non-ablated planners.
BiRRT*PRM*EIT*
Figure 9: Planners with (solid) and without (dashed) shortcutting on the 2D hallway (left) and the 4-arm box-stacking (right) scenarios. All planners are using locally informed sampling. The dashed blue line is the best cost of the non-ablated planners.
@@ V-C2 Shortcutting in the planner
Shortcutting a path after it was found leads to faster convergence than using only informed sampling (Fig. 9) while not affecting the optimality or completeness properties of the planners. This effect is more pronounced in high-dimensional scenarios. If we only do shortcutting, but not locally informed sampling, the planner does not converge to the same solution.
V-C3 Mode sampling strategy
Employing a frontier sampling strategy during the initial search significantly accelerates the discovery of a first solution compared to uniform sampling (Fig. 10) in high dimensional, long horizon scenarios. While not shown here, a greedy sampling strategy gives a further significant speedup for the initial solution but does sacrifice completeness if not done carefully.
BiRRT*PRM*EIT*
Figure 10: Frontier (solid) and uniform (dashed) initial mode sampling strategies on the 2D hallway (left) and 4-arm box-stacking (right) scenarios.
@@ VI Discussion & Limitations
Due to planning in the composite space and treating the problem as a single big planning problem, the planners do not scale well with the number of robots (due to the increase in state-space dimensions) and they do not scale well with the number of goals, as can be seen in the slower convergence in the high dimensional problems. However, we believe that formulating the problem in composite space helps understand how choices such as prioritization affect the continuous space, and this deeper understanding can enable the design of better planners.
We introduce a problem formulation and planners that can deal with all types of task specification. However, we believe that there is a better way to approach the problem than a brute-force search, which the sampling-based planners effectively do here. Despite these limitations, we find that the planning in composite space is likely good enough for many complex industrial settings, where the number of used robots does not go much beyond four robots. Furthermore, with our initial mode sampling strategy defined in Section IV-2, we were able to rapidly reduce the decision space by selecting a feasible mode sequence early in the search, allowing us to find an initial solution quickly.